The equilibrium constant, Kc, is calculated using molar concentrations. For gaseous reactions another form of the equilibrium constant, Kp, is calculated from partial pressures instead of concentrations. These two equilibrium constants are related by the equation

Kp=Kc(RT)?n

where R=0.08206 L?atm/(K?mol), T is the absolute temperature, and ?n is the change in the number of moles of gas (sum moles products - sum moles reactants). For example, consider the reaction

N2(g)+3H2(g)?2NH3(g)

for which ?n=2?(1+3)=?2.

Part A

For the reaction

3A(g)+3B(g)?C(g)

Kc = 40.0 at a temperature of 173?C .

Calculate the value of Kp.

Part B

For the reaction

X(g)+3Y(g)?3Z(g)

Kp = 3.35
Answer #1

A) Kp = Kc ( RT)^delta n

where delta n = no.of moles of gaseous products - no.of moles of gaseous reactants = 1 - ( 3+3) = 1- 6 = -5

R = 0.08206 L atm/k/mole
T = 173 + 273 = 446 K

so Kp = 40.0 X (446 X 0.08206 )^-5 = 40.0 X ( 36.599)^-5 = 74.2 X 1.523 X 10^-8 = 113.0066 X 10^-8

B) delta n = 3 - ( 1 + 3) = 3 -4 = -1

and T = 175 + 273 = 448 K

so 3.35 X 10^2 = Kc ( 0.08206 X 448 )^-1

3.35 X 10^2 = Kc X ( 36.7629)^-1

3.35 X 10^2= Kc X 0.027201

Kc = 3.35 X 10^2 / 0.027201 = 12315.724 (NOTE- if kp = 3.35 X 10^-2, then answer would be 1.2316)